OpenStudy (anonymous):

A force of 5.25 newtons acts on an object of unknown mass at a distance of 6.9..*.108 meters from the center of Earth. To increase the force to 2.5 times its original value, how far should the object be from the center of Earth?

OpenStudy (anonymous):

F1=mMG/d1^2 F2=2.5F1=mMG/d2^2 F2/F1=2.5=d1^2/d2^2 then d2=d1/(2.5)^1/2 d2=6.9x10^8/1.58=4.37x10^8 m hope that helps